OpenStudy (anonymous):

magnitude of 6 root 6-4i root 3

OpenStudy (anonymous):

cant get u

OpenStudy (saifoo.khan):

use the euation editor please.

OpenStudy (saifoo.khan):

equation*

OpenStudy (anonymous):

\[6\sqrt{6}-4i \sqrt{3}\]

OpenStudy (anonymous):

216 + 48

OpenStudy (anonymous):

the magnitude of that

OpenStudy (anonymous):

kanwar, are you on drugs?

OpenStudy (saifoo.khan):

Haha.

OpenStudy (anonymous):

ye, heroin.

OpenStudy (anonymous):

Ah.

OpenStudy (phi):

multiply by the complex conjugate \[ 6\sqrt{6}+4i \sqrt{3} \]and then take the square root of the result

OpenStudy (anonymous):

\[\sqrt{6^2+(4\sqrt{3})^2}\]

OpenStudy (anonymous):

square real part. square imaginary part add take the square root

OpenStudy (anonymous):

\[2\sqrt{21}\]